# Interpretation of generalized Langevin equations

### v3

@@ I.2 The Proposed Equivalence
Our argument is based on the consideration that over any finite time-scale, a theoretical system such as Eq. (3) will have experienced a large enough number of nearly-independent increments that the Generalized Central Limit Theorem should apply [11]. That is, the net increment over any finite time must be drawn from the family of stable distributions, or—if the ”intrinsic” noise represented by the differential update distribution itself has finite variance—a Gaussian distribution in particular [11]. This intuitively dovetails with the more practically-motivated necessary condition that, in the numerical simulation of any continuous-time system, its behavior must not depend sensitively on the simulated timestep; that is, one relatively large step must result in the same distribution (in an ensemble average sense) as the commensurate number of arbitrarily small steps.
@@ I.3 Application to Drag in Turbulent Fluid
(here the constant $c$ sets the time scale, and we set it to $1$ henceforth). When rapid random relative-velocity fluctuations are included, we have:
@@ III Conclusions
We have shown that a class of “nonlinear-stochastic” Langevin equations may be interpreted such that they have well-defined behavior after conversion to an equivalent Itô system. We have applied this theory to a class of nonlinear attracting fixed points to analyze their equilibria via moment relations, and showed that simulations bear out this analysis. This type of equilibrium may be more general than initially apparent, since nearly any isolated attracting fixed point is locally well-approximated by equations of this form.
This conversion technique should lead to more faithful physical modeling, yielding qualitatively different behavior when compared to simplifications which transform a deterministic quantity and add noise afterward. In particular, we have shown that there exists a critical noise level in one such system which leads to divergent moments of its equilibrium, something that cannot occur if $x$-independent noise is simply added after the nonlinear operation. Conversely, our reasoning also leads to the implication that apparent Itô behavior might be driven by any number of nonlinear Langevin processes.

### v4

@@ I.2 The Proposed Equivalence
Our argument is based on the consideration that over any finite time-scale, a theoretical system such as Eq. (3) will have experienced a large enough number of nearly-independent increments that the Generalized Central Limit Theorem should apply [11]. That is, the net increment over any finite time must be drawn from the family of stable distributions, or—if the “intrinsic” noise represented by the differential update distribution itself has finite variance—a Gaussian distribution in particular [11]. This intuitively dovetails with the more practically-motivated necessary condition that, in the numerical simulation of any continuous-time system, its behavior must not depend sensitively on the simulated timestep; that is, one relatively large step must result in the same distribution (in an ensemble average sense) as the commensurate number of arbitrarily small steps.
@@ I.3 Application to Drag in Turbulent Fluid
(here the constant $c$ sets the time scale, and we set it to 111 henceforth). When rapid random relative-velocity fluctuations are included, we have:
@@ III Conclusions
We have shown that a class of “nonlinear-stochastic” Langevin equations may be interpreted such that they have well-defined behavior after conversion to an equivalent Itô system. This type of equilibrium may be more general than initially apparent, since nearly any isolated attracting fixed point is locally well-approximated by equations of this form.
We have applied this theory to cubic and quadratic nonlinear attracting fixed points and arrived at appropriate equivalent Itô systems. These systems notably differ from any naïve system with $x$-independent noise, and Fokker-Planck numerical integration confirms that the implied equilibria of the quadratic case (a model of the velocity of a particle experiencing drag in a turbulent fluid) are quite different. In other work [21] we analyze the equilibria of the cubic system via moment relations and find that its higher moments diverge above a critical value of $\sigma$, and show that simulations bear out this analysis.
This conversion technique should lead to more faithful physical modeling, yielding qualitatively different behavior when compared to simplifications which transform a deterministic quantity and add noise afterward. Conversely, our reasoning also leads to the implication that apparent, observable Itô behavior might be emergent from any number of nonlinear Langevin processes at the theoretical level.